# Never Train from Scratch: Fair Comparison of long-sequence models requires data-driven priors

### v2

@@ 3.3 The Role of Explicit Priors
[…]
@@
[…]
@@
where $\bm{\mathit{1}}$ is the all-ones vector. DLR is significantly simpler to compute compared to S4 and the authors reported it to be as performant as state-of-the-art SSMs on a wide variety of token-level tasks. Hence, it is natural to investigate the conditions under which S4 with its more complex design (eq. 3.3) can be replaced by the simpler DLR. To that end, we evaluate the performance of DLR and S4 on ListOps, Text, Image and PathX tasks as they are the hardest and represent all modalities in LRA. For each model, we experiment with two sets of initializations: (1) random initialization where the state space parameters are initialized from a normal distribution with a small standard deviation, and (2) “structured” initialization recommended by the respective authors aimed at capturing long-range dependencies.
@@ C.1.2 Comparing S4 and Transformers
Transformer + Masked ${}^{\dagger}$ 512 6 8 1024 Mean 1e-4,1e-3 64,128 0.1 CE
Transformer + Causal ${}^{\dagger}$ 512 6 8 1024 Last 5e-4, 5e-4 128,32 0.1 CE
@@
Transformer + Masked ${}^{\dagger}$ 512 6 8 1024 Mean 1e-4,5e-4 64,32 0.1 CE
Transformer + Causal ${}^{\dagger}$ 512 6 8 1024 Mean 1e-4,1e-3 64,32 0.1 CE
@@
Transformer + Masked ${}^{\dagger}$ 128 4 4 512 Mean 5e-4,5e-3 16,32 0 CE
Transformer + Causal ${}^{\dagger}$ 128 4 4 512 Mean 1e-3,5e-3 48,32 0 CE
@@
Transformer + Masked ${}^{\dagger}$ 64 3 4 128 Max 1e-3,1e03 16,32 0 L1
Transformer + Causal ${}^{\dagger}$ 64 3 4 128 Max 1e-3,5e-3 32,32 0 L1
@@
Transformer + Masked ${}^{\dagger}$ 128 4 8 128 Mean 5e-4,1e-3 16,16 0 CE
Transformer + Masked ${}^{\dagger}$ 128 4 8 128 Mean 1e-3,1e-3 256,128 0 CE
@@
Transformer + Masked ${}^{\dagger}$ 128 4 8 128 Max 5e-4,5e-4 32,8 0 CE
Transformer + Masked ${}^{\dagger}$ 128 4 8 128 Max 5e-4,1e-3 256,8 0 CE

### v4

@@ 3.3 The Role of Explicit Priors
where $\bm{\mathit{\Lambda}},\bm{\mathit{P}},\bm{\mathit{Q}},\bm{\mathit{B}},\bm{\mathit{C}},\Delta$ are learnable parameters and $\bm{\mathit{\Lambda}}\in\textit{Diag}(\mathbb{C}^{N\times N}),\bm{\mathit{P}},\bm{\mathit{Q}}\in\mathbb{C}^{N\times 1}$. In addition to this parameterization, S4 uses a principled initialization method aimed towards a slow decay of the kernel (w.r.t. timestep $k$) in order to facilitate capturing long-range dependencies.
@@
$\vec{\bm{x}}_{n}=\bm{\mathit{\Lambda}}\vec{\bm{x}}_{n-1}+\bm{\mathit{1}}u_{n}\quad\bm{\mathit{\Lambda}}\in\textit{diag}(\mathbb{C}^{N\times N})$ (3)
@@
where $\bm{\mathit{1}}$ is the all-ones vector. DLR is significantly simpler to compute compared to S4 and the authors reported it to be as performant as state-of-the-art SSMs on a wide variety of token-level tasks. Hence, it is natural to investigate the conditions under which S4 with its more complex design (eq. 2) can be replaced by the simpler DLR. To that end, we evaluate the performance of DLR and S4 on ListOps, Text, Image and PathX tasks as they are the hardest and represent all modalities in LRA. For each model, we experiment with two sets of initializations: (1) random initialization where the state space parameters are initialized from a normal distribution with a small standard deviation, and (2) “structured” initialization recommended by the respective authors aimed at capturing long-range dependencies.
@@ 3.7 Additional Experiments
4 Acknowledgments
We thank Amir Globerson for insightful discussions and his support throughout the course of this work. This research was partially supported by Facebook and by the Yandex Initiative for Machine Learning, and the European Research Council (ERC) under the European Union Horizons 2020 research and innovation programme (grant ERC DELPHI 802800).
@@ C.1.2 Comparing S4 and Transformers
Transformer + Masked SPT 512 6 8 1024 Mean 1e-4,1e-3 64,128 0.1 CE
Transformer + Causal SPT 512 6 8 1024 Last 5e-4, 5e-4 128,32 0.1 CE
@@
Transformer + Masked SPT 512 6 8 1024 Mean 1e-4,5e-4 64,32 0.1 CE
Transformer + Causal SPT 512 6 8 1024 Mean 1e-4,1e-3 64,32 0.1 CE
@@
Transformer + Masked SPT 128 4 4 512 Mean 5e-4,5e-3 16,32 0 CE
Transformer + Causal SPT 128 4 4 512 Mean 1e-3,5e-3 48,32 0 CE
@@
Transformer + Masked SPT 64 3 4 128 Max 1e-3,1e03 16,32 0 L1
Transformer + Causal SPT 64 3 4 128 Max 1e-3,5e-3 32,32 0 L1
@@
Transformer + Masked SPT 128 4 8 128 Mean 5e-4,1e-3 16,16 0 CE
Transformer + Masked SPT 128 4 8 128 Mean 1e-3,1e-3 256,128 0 CE
@@
Transformer + Masked SPT 128 4 8 128 Max 5e-4,5e-4 32,8 0 CE
Transformer + Masked SPT 128 4 8 128 Max 5e-4,1e-3 256,8 0 CE